Given f=50Hz, determine how long it will take the sinusoidal waveform to pass through an angle of 50o


1
Expert's answer
2021-09-02T10:42:41-0400

f= 50 Hz

"T = \frac{1}{f}= \frac{1}{50}= 0.02 \;s"

One sinusoidal wave completes 0.02 s in 360°.

Proportion:

360° - 0.02 s

50 ° - x

"x = \frac{0.02 \times 50}{360}= 0.00277 \;s = 2.77 \;msec"

Answer: 2.77 msec